# Statistical Inference for Temporal Difference Learning with Linear Function Approximation

### v1

@@ B.1 High-probability convergence guarantee for the original TD estimation error
$\widetilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (75)
B.3 Proof of Theorem 3.2
@@
For the last term on the right-hand-side of (B.3), recall from the definitions of $\bm{\omega}_{j}$ (78) and $\bm{\omega}_{j}^{(i)}$ (83), for any $j>i$,
@@
since $\mathbb{E}_{j^{\prime}-1}[\bm{A}_{j^{\prime}}-\bm{A}]=\bm{0}$. This effectively means that the last term on the right-hand-side of (B.3) is 00. Therefore, the difference bewteen $\Delta$ and $\Delta^{(i)}$ is bounded by
@@
By combining (B.3), (B.3),(B.3) with (86b), we obtain
@@
With $I_{1}$ bounded by (B.3), the proof now boils down to bounding $I_{2}$. Notice that (86a) implies
@@
A direct combination of (89), (90) and (B.3) yields the following bound for the third term of (79):
@@
$\left\|\bm{\Lambda}^{\star-\frac{1}{2}}\overline{\bm{\Lambda}}_{T}\bm{\Lambda}^{\star-\frac{1}{2}}-\bm{I}\right\|_{\mathsf{F}}=\left\|\bm{\Lambda}^{\star-\frac{1}{2}}(\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{\Lambda}^{\star-\frac{1}{2}}\right\|_{\mathsf{F}}=O(T^{\alpha-1}).$
Consequently, the $O(T^{\alpha-1})$ rate by which $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ converges to $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ cannot be further improved.
B.5 Proof of Theorem 3.4
Applying the same reasoning as in the proof of Theorem 3.3,
@@
$\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}^{-\top}-\bm{A}\bm{\Gamma}\bm{A}$
$=\overline{\bm{A}}^{-1}(\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma})\overline{\bm{A}}^{-\top}+\left[\overline{\bm{A}}^{-1}\bm{\Gamma}\overline{\bm{A}}^{-\top}-\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top}\right].$ (93)
@@
Let $\overline{\bm{A}}$ be defined as in (20) with $\{(s_{t},s_{t}^{\prime})\}_{0\leq t\leq T}$ being $i.i.d.$ samples. It can then be guaranteed with probability at least $1-\delta$ that
$\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq\frac{4\sqrt{2}}{\lambda_{0}(1-\gamma)\sqrt{T}}\sqrt{\log\frac{3}{\delta}},$
@@
Based on Lemmas B.4 and B.5, we now consider the estimation error of $\widehat{\bm{\Lambda}}_{T}$. By triangle inequality, (B.5) implies that
$\left\|\bm{A}(\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\right\|_{\mathsf{F}}\leq\underset{I_{1}}{\underbrace{\|\overline{\bm{A}}^{-1}\bm{A}\|^{2}\|\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}\|_{\mathsf{F}}}}+\underset{I_{2}}{\underbrace{\left\|\bm{A}\overline{\bm{A}}^{-1}\bm{\Gamma}\overline{\bm{A}}^{-\top}\bm{A}^{\top}-\bm{\Gamma}\right\|_{\mathsf{F}}}}.$
@@
$I_{2}$ $\leq\|(\bm{A}\overline{\bm{A}}^{-1}-\bm{I})\bm{\Gamma}\|_{\mathsf{F}}+\|\bm{\Gamma}(\overline{\bm{A}}^{-\top}\bm{A}^{\top}-\bm{I})\|_{\mathsf{F}}+\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{\Gamma}(\overline{\bm{A}}^{-\top}-\bm{A}^{-\top})\|_{\mathsf{F}}$
$\leq 2\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}+\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}^{2},$
@@
Here, $I_{1}$ and $I_{2}$ measures the difference between the non-asymptotic and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$, and are bounded by Theorems 3.2 and 3.3 respectively. Specifically,
@@ Appendix C Proof of supportive lemmas and propositions
holds uniformly for all $0\leq t\leq T$. Hence the vector Azuma’s inequality A.16 directly implies that
$\|\overline{\bm{A}}-\bm{A}\|_{\mathsf{F}}$ $=\left\|\mathbf{vec}(\overline{\bm{A}}-\bm{A})\right\|\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}$
with probability at least $1-\delta$. In what follows, we aim to transform this error bound to the error bound for $\overline{\bm{A}}^{-1}$. Towards this end, we observe that condition 94 guarantees with probability at least $1-\delta$,
$\left\|\bm{A}^{-1}(\overline{\bm{A}}-\bm{A})\right\|$ $\leq\|\bm{A}^{-1}\|\|\overline{\bm{A}}-\bm{A}\|$
@@
$\|(\overline{\bm{A}}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq 2\|(\overline{\bm{A}}-\bm{A})\bm{A}^{-1}\|_{\mathsf{F}}$ $\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}\|\bm{A}^{-1}\|$
@@
$\mathbb{E}\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ $=\frac{1}{T}\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})+\widetilde{C}T^{-\alpha-1}=\frac{1}{T}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(T^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1}),$
@@
$\quad\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
$\quad\leq\frac{2}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
$\quad\leq\frac{2}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
$\quad\leq\frac{4}{T}\sum_{t=1}^{T}4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}$
$\quad=4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}.$ (105)

### v3

@@ B.1.1 High-probability convergence guarantee for the original TD estimation error
$\tilde{C}$ $=\eta_{0}^{2}\left(\frac{2}{\beta}\right)^{\frac{2+\alpha}{2(1-\alpha)}}(6\|\bm{\theta}^{\star}\|_{2}+1)\sqrt{\frac{1}{(1-\alpha)(2\alpha-1)}}.$ (78)
B.1.3 The necessity of the condition (71)
An important fact to note about Theorem 3.1 is that it relies on choosing an appropriate initial stepsize $\eta_{0}$ dependent of $\delta$. The relationship is specified in Equation (71). This constraint, albeit only related to $\delta$ in logarithm terms, prevents $\delta$ to converge to 00 for any fixed $\eta_{0}$. One might naturally wonder if such a constraint could be relaxed and the convergence of $\overline{\bm{\Delta}}_{T}$ can be guaranteed with high probability for any fixed $\eta_{0}$ and arbitrarily small $\delta$. The following proposition shows that the answer is negative, so some form of constraint on $\delta$ is necessary for a Chernoff-style convergence guarantee.
Proposition B.2.
Consider the TD algorithm (7) with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, where $\eta_{0}$ is fixed and $\alpha\in(\frac{1}{2},1)$. There exists an MDP, such that for any bi-variate polynomial function $\mathsf{poly}(\cdot,\cdot)$, there exists $\delta\in(0,1)$ and $T\in\mathbb{N}$, such that
$\mathbb{P}\left(\|\overline{\bm{\Delta}}_{T}\|_{2}\geq\mathsf{poly}\left(T,\log\frac{1}{\delta}\right)\right)\geq\delta.$
Proof: Consider the following MDP, where $|\mathcal{S}|=3$, $\gamma=0.9$, while the transition matrix $\bm{P}$ and the reward vector $\bm{r}$ are
$\bm{P}=\begin{pmatrix}0.1&0.1&0.8\\ 0.1&0.1&0.8\\ 0.1&0.1&0.8\end{pmatrix},\quad\text{and}\quad\bm{r}=\begin{pmatrix}0.1\\ 0.1\\ 1\end{pmatrix}.$
[…]
$\bm{A}=0.54475,\quad\bm{b}=0.785,\quad\text{and}\quad\bm{\theta}^{\star}=1.441.$
For any $T\in\mathbb{N}$, use $\mathcal{E}_{T}$ to denote the event that $s_{t}=1$ and $s_{t}^{\prime}=2$ for all $t\in[T]$. In the case of $i.i.d.$ samples, clearly $\mathbb{P}(\mathcal{E}_{T})=0.01^{T}$. For any bi-variate polynomial function $\mathsf{poly}(\cdot,\cdot)$, let $\delta=0.01^{T}$, then $\log\frac{1}{\delta}=T\log 100$. Therefore, $\mathsf{poly}(T,\log\frac{1}{\delta})=\mathsf{poly}(T,T\log 100)$ reduces to a univariate polynomial function of $T$. It suffices to show that under the event $\mathcal{E}_{T}$, the norm of $\overline{\bm{\Delta}}_{T}$ grows faster than any polynomial function of $T$.
Under the event $\mathcal{E}_{T}$, direct calculation yields $\bm{A}_{t}=-0.2$ and $\bm{b}_{t}=-0.05$. Hence, the estimation error $\bm{\Delta}_{t}$ is featured by the iterative relation
$\bm{\Delta}_{t}$ $=\bm{\Delta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\Delta}_{t-1}+\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})$
$=(1+0.2\eta_{t})\bm{\Delta}_{t-1}+(0.2\bm{\theta}^{\star}-0.05)\eta_{t}.$
It is easy to show by induction that, when $\bm{\theta}_{0}=0$ and hence $\bm{\Delta}_{0}=-\bm{\theta}^{\star}$,
$\bm{\Delta}_{t}$ $=(0.2\bm{\theta}^{\star}-0.05)\sum_{i=1}^{t}\eta_{i}\prod_{k=i+1}^{t}(1+0.2\eta_{k})+\prod_{k=1}^{t}(1+0.2\eta_{k})\bm{\Delta}_{0}$
$=5(0.2\bm{\theta}^{\star}-0.05)\sum_{i=1}^{t}\left(\prod_{k=i}^{t}-\prod_{k=i+1}^{t}\right)(1+0.2\eta_{k})+\prod_{k=1}^{t}(1+0.2\eta_{k})(-\bm{\theta}^{\star})$
$=5(0.2\bm{\theta}^{\star}-0.05)\left(\prod_{k=1}^{t}(1+0.2\eta_{k})-1\right)-\prod_{k=1}^{t}(1+0.2\eta_{k})\bm{\theta}^{\star}$
$=-0.25\prod_{k=1}^{t}(1+0.2\eta_{k})+(0.25-\bm{\theta}^{\star})$
$=-0.25\prod_{k=1}^{t}(1+0.2\eta_{k})-1.191.$
Consequently, the averaged TD estimation error is featured by
$\overline{\bm{\Delta}}_{T}=-1.191-0.25\frac{1}{T}\sum_{t=1}^{T}\prod_{k=1}^{t}(1+0.2\eta_{k}),$
and therefore
$\left\|\overline{\bm{\Delta}}_{T}\right\|_{2}=1.191+0.25\frac{1}{T}\sum_{t=1}^{T}\prod_{k=1}^{t}(1+0.2\eta_{k})$
We now consider the growth of $\left\|\overline{\bm{\Delta}}_{T}\right\|_{2}$ with polynomial-decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, where $\alpha\in(\frac{1}{2},1)$. In this scenario, the norm of $\overline{\bm{\Delta}}_{T}$ is bounded by
$\left\|\overline{\bm{\Delta}}_{T}\right\|_{2}$ $=1.191+0.25\frac{1}{T}\sum_{t=1}^{T}\prod_{k=1}^{t}(1+0.2\eta_{k})$
$=1.191+\frac{0.25}{T}\cdot\sum_{t=1}^{T}\frac{1}{0.2\eta_{t}}\left(\prod_{k=1}^{t}-\prod_{k=1}^{t-1}\right)(1+0.2\eta_{k})$
$>1.191+\frac{0.25}{0.2\eta_{T}T}\sum_{t=1}^{T}\left(\prod_{k=1}^{t}-\prod_{k=1}^{t-1}\right)(1+0.2\eta_{k})$
$=1.191+\frac{5}{4\eta_{0}}T^{-\alpha-1}\left(\prod_{k=1}^{T}(1+0.2\eta_{0}k^{-\alpha})-1\right).$
Here, since $k^{-\alpha}\leq 1$, it can be guaranteed that
$\prod_{k=1}^{T}(1+0.2\eta_{0}k^{-\alpha})$ $>\prod_{k=1}^{T}(1+0.2\eta_{0})^{k^{-\alpha}}$
$=\exp\left(\log(1+0.2\eta_{0})\sum_{k=1}^{T}k^{-\alpha}\right)$
$>\exp\left(\log(1+0.2\eta_{0})\cdot\frac{T^{1-\alpha}}{1-\alpha}\right).$
Take $u=T^{1-\alpha}$, then the norm of $\overline{\bm{\Delta}}_{T}$ is bounded below by
$\left\|\overline{\bm{\Delta}}_{T}\right\|_{2}>1.191+\frac{5}{4\eta_{0}}u^{-\frac{\alpha+1}{1-\alpha}}\left[\exp\left(\frac{\log(1+0.2\eta_{0})}{1-\alpha}u\right)-1\right],$
which grows exponentially with $u$ and hence faster than any polynomial function of $T$. This completes the proof of the proposition. ∎
B.2 Proof of Theorem 3.2
By triangle inequality, we firstly decompose the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and its asymptotic distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ by
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))+d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})).$ (79)
Here, recall that $\overline{\bm{\Lambda}}_{T}$ is defined as in (49). The following two theorems bound the two terms on the right-hand-side of (79) respectively; Theorem 3.2 follows directly by combining them with (79).
Theorem B.3.
Consider the TD with Polyak-Ruppert averaging with independent samples and decaying stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$ for $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}<1/2$. Then,
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\tilde{C}\;T^{-\frac{1}{2}},$
where $\tilde{C}$ is a problem related quantity independent of $T$.
Theorem B.4.
Let $\bm{\Lambda}^{\star}$ and $\overline{\bm{\Lambda}}_{T}$ be defined as in (12) and (49), respectively, with the stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, for all $t>0$, any $\alpha\in(\frac{1}{2},1)$ and $\eta_{0}<1/2$. Then,
$d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\leq d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\tilde{C}^{\prime}\;T^{2\alpha-2},$
where $\tilde{C}^{\prime}$ is a problem related quantity independent of $T$.
B.2.1 Proof of Theorem B.3
@@ B.1.3 The necessity of the condition (71)
For the last term on the right-hand-side of (B.2.1), recall from the definitions of $\bm{\omega}_{j}$ (82) and $\bm{\omega}_{j}^{(i)}$ (87), for any $j>i$,
@@
since $\mathbb{E}_{j^{\prime}-1}[\bm{A}_{j^{\prime}}-\bm{A}]=\bm{0}$. This effectively means that the last term on the right-hand-side of (B.2.1) is 00. Therefore, the difference bewteen $\Delta$ and $\Delta^{(i)}$ is bounded by
@@
By combining (B.2.1), (B.2.1),(B.2.1) with (90b), we obtain
@@
With $I_{1}$ bounded by (B.2.1), the proof now boils down to bounding $I_{2}$. Notice that (90a) implies
@@
A direct combination of (93), (94) and (B.2.1) yields the following bound for the third term of (83):
@@
$d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\approx\left\|\bm{\Lambda}^{\star-\frac{1}{2}}\overline{\bm{\Lambda}}_{T}\bm{\Lambda}^{\star-\frac{1}{2}}-\bm{I}\right\|_{\mathsf{F}}$
$=\left\|\bm{\Lambda}^{\star-\frac{1}{2}}(\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{\Lambda}^{\star-\frac{1}{2}}\right\|_{\mathsf{F}}=O(T^{\alpha-1}).$
Consequently, the $O(T^{\alpha-1})$ rate by which $\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T})$ converges to $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$ cannot be further improved. As a direct consequence, when $\alpha\in(\frac{2}{3},1)$, since $\alpha-1>-\frac{1}{3}>-\frac{\alpha}{2}$, it can be guaranteed through triangle inequality that
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\geq d_{\mathsf{C}}(\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))-d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\overline{\bm{\Lambda}}_{T}))$
$\geq\Theta(T^{\alpha-1})-O(T^{-\frac{\alpha}{2}})\geq\Theta(T^{-\frac{1}{3}}).$
B.3 Proof of Theorem 3.3
Applying the same reasoning as in the proof of Theorem B.4,
@@
$\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}$ $=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}\overline{\bm{A}}_{T}^{-\top}-\bm{A}\bm{\Gamma}\bm{A}$
$=\overline{\bm{A}}_{T}^{-1}(\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma})\overline{\bm{A}}_{T}^{-\top}+\left[\overline{\bm{A}}_{T}^{-1}\bm{\Gamma}\overline{\bm{A}}_{T}^{-\top}-\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top}\right].$ (97)
@@
Let $\overline{\bm{A}}_{T}$ be defined as in (14) with $\{(s_{t},s_{t}^{\prime})\}_{0\leq t\leq T}$ being $i.i.d.$ samples. It can then be guaranteed with probability at least $1-\delta$ that
$\|(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq\frac{4\sqrt{2}}{\lambda_{0}(1-\gamma)\sqrt{T}}\sqrt{\log\frac{3}{\delta}},$
@@
Based on Lemmas B.7 and B.8, we now consider the estimation error of $\widehat{\bm{\Lambda}}_{T}$. By triangle inequality, (B.3) implies that
$\left\|\bm{A}(\widehat{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star})\bm{A}^{\top}\right\|_{\mathsf{F}}\leq\underset{I_{1}}{\underbrace{\|\overline{\bm{A}}_{T}^{-1}\bm{A}\|^{2}\|\widehat{\bm{\Gamma}}_{T}-\bm{\Gamma}\|_{\mathsf{F}}}}+\underset{I_{2}}{\underbrace{\left\|\bm{A}\overline{\bm{A}}_{T}^{-1}\bm{\Gamma}\overline{\bm{A}}_{T}^{-\top}\bm{A}^{\top}-\bm{\Gamma}\right\|_{\mathsf{F}}}}.$
@@
$I_{2}$ $\leq\|(\bm{A}\overline{\bm{A}}_{T}^{-1}-\bm{I})\bm{\Gamma}\|_{\mathsf{F}}+\|\bm{\Gamma}(\overline{\bm{A}}_{T}^{-\top}\bm{A}^{\top}-\bm{I})\|_{\mathsf{F}}+\|(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\bm{\Gamma}(\overline{\bm{A}}_{T}^{-\top}-\bm{A}^{-\top})\|_{\mathsf{F}}$
$\leq 2\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}+\|\bm{\Gamma}\|\|\bm{A}(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\|_{\mathsf{F}}^{2},$
@@
Here, $I_{1}$ and $I_{2}$ measures the difference between the non-asymptotic and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$, and are bounded by Theorems B.3 and B.4 respectively. Specifically,
@@ Appendix C Proof of supportive lemmas and propositions
holds uniformly for all $0\leq t\leq T$. Hence the vector Azuma’s inequality A.18 directly implies that
$\|\overline{\bm{A}}_{T}-\bm{A}\|_{\mathsf{F}}$ $=\left\|\mathbf{vec}(\overline{\bm{A}}_{T}-\bm{A})\right\|\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}$
with probability at least $1-\delta$. In what follows, we aim to transform this error bound to the error bound for $\overline{\bm{A}}_{T}^{-1}$. Towards this end, we observe that condition (98) guarantees with probability at least $1-\delta$,
$\left\|\bm{A}^{-1}(\overline{\bm{A}}_{T}-\bm{A})\right\|$ $\leq\|\bm{A}^{-1}\|\|\overline{\bm{A}}_{T}-\bm{A}\|$
@@
$\|(\overline{\bm{A}}_{T}^{-1}-\bm{A}^{-1})\bm{A}\|_{\mathsf{F}}\leq 2\|(\overline{\bm{A}}_{T}-\bm{A})\bm{A}^{-1}\|_{\mathsf{F}}$ $\leq\frac{4\sqrt{2}}{\sqrt{T}}\sqrt{\log\frac{3}{\delta}}\|\bm{A}^{-1}\|$
@@
$\mathbb{E}\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$ $=\frac{1}{T}\mathsf{Tr}(\overline{\bm{\Lambda}}_{T})+\tilde{C}T^{-\alpha-1}=\frac{1}{T}\mathsf{Tr}(\bm{\Lambda}^{\star})+o(T^{-1})\leq\frac{\mathsf{Tr}(\bm{\Gamma})}{\lambda_{0}^{2}(1-\gamma)^{2}T}+o(T^{-1}),$
@@
$\left\|\frac{1}{T}\sum_{t=1}^{T}\left[(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right]\right\|_{\mathsf{F}}$ (109)
$\leq\frac{1}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}+(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}\right\|_{\mathsf{F}}$
$\leq\frac{2}{T}\sum_{t=1}^{T}\left\|(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right\|_{\mathsf{F}}$
$\leq\frac{2}{T}\sum_{t=1}^{T}\sqrt{\mathsf{Tr}\left((\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\Delta}}_{T})^{\top}(\bm{A}_{t}\overline{\bm{\Delta}}_{T})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}\right)}$
$\leq\frac{4}{T}\sum_{t=1}^{T}4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}$
$=4(2\|\bm{\theta}^{\star}\|_{2}+1)\|\overline{\bm{\Delta}}_{T}\|_{2}.$ (110)